OpenStudy (anonymous):

I am currently converting Rectangular points into Polar points. The point is (3,-1) I have so far converted this much: X= 3 Y=-1 T=Theta tan T= Y/X T=arctan(-3) T=? T=? x^2+y^2=r^2 r=+or- sqrt 10

OpenStudy (anonymous):

\[\tan \theta = \left(\begin{matrix}x \\ y\end{matrix}\right)\] \[\theta = \arctan (-3)\] \[\theta = ?\] \[\theta = ?\]

OpenStudy (dumbcow):



OpenStudy (anonymous):

I meant \[\left(\begin{matrix}y \\ x\end{matrix}\right)\] sorry. This might go under Unit Circle Pre-Calc stuff.

OpenStudy (dumbcow):

right so the polar coordinates of point (3,-1) is (sqrt10 , arctan(-1/3))

OpenStudy (anonymous):

would there be a specific coordinate on the unit circle?, such as \[\left(\begin{matrix}\pi \\ 3\end{matrix}\right)\] or would the general statement of arctan be good? because I know arctan(-1) \[\theta = \left(\begin{matrix}\pi \\ 4\end{matrix}\right)\] and \[\theta = \left(\begin{matrix}3\pi \\ 4\end{matrix}\right)\]

OpenStudy (dumbcow):

no not for -1/3 , the base angles pi/6 , pi/4 , pi/3 have tangents of 1/sqrt3 , 1 , sqrt3 you can only approximate angle with calculator